### v1

@@ 3.3 Results: No Covariates
The simulation study with equal sample sizes demonstrates that the UN–DID estimator is unbiased. To assess unbiasedness, we proceed to generate kernel density plots for both conventional and UN–DID estimates of the ATT on a shared axis. The kernel densities for unequal sample sizes are shown in Figure 2. In Panel (a), the kernel density estimates for Case 1 are presented, and in Panel (b), the kernel density estimates for Case 2 are presented. Notably, for Case 1, both kernels are centered around 0, aligning with the true value of the ATT. Similarly, for Case 2, both kernels are centered around 0.1, which is the true value of the ATT for Case 2. This finding demonstrates that both conventional and UN–DID estimation methods exhibit unbiasedness, as evidenced by the distributions of estimates being centered around the true values of the ATT. A similar pattern is found with equal sample sizes, in a figure not included.
In order to prove whether the UN–DID approaches the true value as sample size increases, we repeat the Monte Carlo simulation by increasing the sample size to 1,000, 2,000, 4,000, 8,000, 10,000, and 50,000 while maintaining equal sample sizes between both silos. Subsequently, we compute the mean squared error (MSE) between the estimated ATT using UN–DID and the true ATT from the underlying DGP according to Equation (26):
$MSE(\widehat{ATT})=\frac{1}{1000}\sum_{j}^{1000}\biggl{(}\widehat{ATT_{j}}^{Unpooled}-ATT_{0}\biggr{)}^{2}.$ (26)
@@
$MSE(\widehat{SE})=\frac{1}{1000}\sum_{j}^{1000}\biggl{(}\widehat{SE}(\widehat{ATT_{j}})^{Unpooled}-SE_{0}\biggr{)}^{2}.$ (28)
The findings are shown in Figure 3, where the y-axis represents the mean squared error (MSE) and the x-axis corresponds to the sample size. The graph illustrates a decreasing trend in MSE for both the estimated ATT and the standard error as the sample size increases. The results for unequal sample sizes are illustrated in Figure 4. We observe that the MSE for the case with unequal sample sizes is higher than for the case with equal sample sizes but still converges to 0 as sample size increases.
Figure 3: No Covariates with Equal Sample Sizes: Mean Squared Errors
Figure 4: No Covariates with Equal Sample Sizes: Mean Squared Errors
4 UN–DID When Strong Parallel Trends Do Not Hold
In the previous section, we assumed that strong parallel trends is plausible between the treated and the control silos. In this section, we relax the assumption so that parallel trends is plausible only conditional on covariates. In other words, we relax Assumption 2 and introduce the conditional parallel trends assumption.
@@ 4 UN–DID When Strong Parallel Trends Do Not Hold
$\begin{split}\biggr{[}E[Y_{i}(0)|D_{i}=1,t=1,X_{i,t}]-E[Y_{i}(0)|D_{i}=1,t=0,X_{i,t}]\biggr{]}\\ =\biggr{[}E[Y_{i}(0)|D_{i}=0,t=1,X_{i,t}]-E[Y_{i}(0)|D_{i}=0,t=0,X_{i,t}]\biggr{]}.\end{split}$
Conditional parallel trends assumption states that the evolution of outcomes between the treated and control groups are the same in the pre-intervention period, conditional on covariates. Under the conditional parallel trends assumption, we can also switch the ATT, shown in Equation (3), to the ATT conditional on covariates:
@@
With covariates, we can extend Assumption 6 such that matrices of the covariates, $X_{T}$ and $X_{C}$ for the treated and the untreated groups, respectively, can be stacked together into a common matrix $X$:
@@
This stacking is not feasible when data are unpoolable. With unpoolable data, when we gain access to the data from the treated silo, we do not have any information on the data from the control group, or both $Y_{C}$ and $X_{C}$. Similarly, when we gain access to the data from the untreated silo, we do not have any information on the data for the treated silo, or both $Y_{T}$ and $X_{T}$. Because we do not observe the covariates from other groups, we are unable to use non-parametric methods such as the inverse probability weighting (IPW), outcome regression (OR), and the doubly robust methods (DR–DID) to estimate the ATT (Abadie, 2005; Sant’Anna and Zhao, 2020; Heckman et al., 1997). Researchers typically incorporate covariates into a model to ensure the plausibility of parallel trends given the covariates, particularly when strong parallel trends do not hold (Heckman et al., 1997; Abadie, 2005). In our paper, we are also including covariates that are only required for parallel trends to be plausible, conditional on these covariates. The literature on DID with covariates focused on scenarios where covariates are entirely time-invariant (Abadie, 2005; Callaway and Sant’Anna, 2021; Caetano et al., 2022; Roth et al., 2022; Sant’Anna and Zhao, 2020). When dealing with time-varying covariates, researchers in the literature frequently opt to select a specific value of these covariates during a pre-treatment period, treating it as if it were a time-invariant covariate (Caetano et al., 2022). For valid estimation of the ATT with covariates, we require that the covariates are unaffected by the treatment on the basis of the conditional independence assumption.
@@
We will begin by examining the simplest case, where we introduce a single covariate into both the conventional and UN–DID regressions. With a single covariate, the conventional regression is expressed as
@@
Likewise, we adjust the UN–DID regression to incorporate covariates:
@@
Including time-invariant covariates does not complicate DID analysis, as demonstrated in the following subsection. However, with time-varying covariates, Caetano et al. (2022) has demonstrated that the two-way fixed effects (TWFE) DID regression can provide biased estimate of the ATT due to negative weighting and weight-reversal issues. The problem is more pronounced when the slope parameters of the covariate are uncommon between silos. In this paper, we recognize the complexity introduced by the heterogeneity in data distribution between silos, which can present challenges when analyzing siloed data (Li et al., 2022). This challenge arises from the potential uncommon slope parameters of covariates between silos. In our current analysis, we do not directly address this and impose the following assumption:
Assumption 9: Common Slope Parameters of Covariates Between Silos
$\gamma_{2}^{T^{0}}=\gamma_{2}^{C^{0}}.$ (34)
The common slope parameters of covariate between silos states that the slope or the true coefficient of the covariates is the same for both the treated and untreated silos. Therefore, $\beta_{4}^{0}=\beta_{4}^{T^{0}}=\beta_{4}^{C^{0}}$.
4.1 Equivalence of the Estimated ATT Between Conventional and UN–DID Methods with Covariates
Conventional Regression
For the conventional regression, we assume that Assumption 1 and assumptions 3 to 7 hold. With a single covariate, we can expand on the conventional regression in the following way:
In Equation (35), $\hat{\beta_{3}}$ is the parameter of interest, which is the estimate of the ATT conditional on covariates, and is the sample analogue of Equation (29). According to the FWL theorem, $\hat{\beta_{3}}$ will be the same as the coefficient of $(P_{t}*D_{i}-\widehat{P_{t}*D_{i}})$ from the regression shown in Equation (37). Here, $(P_{t}*D_{i}-\widehat{P_{t}*D_{i}})$ are the residuals from Equation (36):
@@ 4.1 Equivalence of the Estimated ATT Between Conventional and UN–DID Methods with Covariates
From Equation (36), we calculate residuals ($P_{t}*D_{i}-\widehat{P_{t}*D_{i}}$) for each observation based on their treatment status and time period. For treated observations in the pre-intervention period, the residual is given by $-\hat{\alpha_{0}}-\hat{\alpha_{2}}-\hat{\alpha_{3}}X_{i,t}$, while for the post-intervention period, it is $1-\hat{\alpha_{0}}-\hat{\alpha_{1}}-\hat{\alpha_{2}}-\hat{\alpha_{3}}X_{i,t}$. Similarly, untreated observations have residuals of $-\hat{\alpha_{0}}-\hat{\alpha_{3}}X_{i,t}$ in the pre-intervention period and $-\hat{\alpha_{0}}-\hat{\alpha_{1}}-\hat{\alpha_{3}}X_{i,t}$ in the post-intervention period. Subsequently, $\widehat{\beta_{3}}$ from Equation (37) can be estimated using the following OLS formula:
$\begin{split}\hat{\beta}_{3}=&\sum_{i}\sum_{t}\frac{Y_{i,t}\left(P_{t}*D_{i}-P_{t}*D_{i}\right)}{\left(P_{t}*D_{i}-P_{t}*D_{i}\right)^{2}}\end{split}$ (38)
@@
$\begin{split}&\sum_{i\in t=0,T}Y_{i,0}^{T}\left(-\hat{\alpha}_{0}-\hat{\alpha}_{2}-\hat{\alpha_{3}}X_{i,0}\right)+\sum_{i\in t=1,T}Y_{i,1}^{T}\left(1-\hat{\alpha}_{0}-\hat{\alpha}_{1}-\hat{\alpha}_{2}-\hat{\alpha_{3}}X_{i,1}\right)+\\ &\sum_{i\in t=0,T}Y_{i,1}^{T}\left(1-\hat{\alpha}_{0}-\hat{\alpha}_{1}-\hat{\alpha}_{2}-\hat{\alpha_{3}}X_{i,0}\right)+\sum_{i\in t=1,C}Y^{C}_{i,1}\left(-\hat{\alpha}_{0}-\hat{\alpha}_{1}-\hat{\alpha_{3}}X_{i,1}\right).\end{split}$
@@
$\begin{split}G=&\sum_{i\in t=0,T}\left(-\hat{\alpha}_{0}-\hat{\alpha}_{2}-\hat{\alpha_{3}}X_{i,t}\right)^{2}+\sum_{i\in t=1,T}\left(1-\hat{\alpha}_{0}-\hat{\alpha}_{1}-\hat{\alpha}_{2}-\hat{\alpha_{3}}X_{i,t}\right)^{2}\\ &+\sum_{i\in t=0,C}\left(-\hat{\alpha}_{0}-\hat{\alpha_{3}}X_{i,t}\right)^{2}+\sum_{i\in t=1,C}\left(-\hat{\alpha}_{0}-\hat{\alpha}_{1}-\hat{\alpha_{3}}X_{i,t}\right)^{2}.\end{split}$
Combining the numerator and the denominator, $\hat{\beta_{3}}$ can be written as
$\begin{split}\hat{\beta}_{3}=&\frac{\sum_{i\in t=1,T}Y_{i,1}^{T}\left(1-\hat{\alpha}_{0}-\hat{\alpha}_{1}-\hat{\alpha}_{2}-\hat{\alpha_{3}}X_{i,1}\right)}{G}-\frac{\sum_{i\in t=0,T}Y_{i,0}^{T}\left(\hat{\alpha}_{0}+\hat{\alpha}_{2}+\hat{\alpha_{3}}X_{i,0}\right)}{G}\\ &-\frac{\sum_{i\in t=1,C}Y_{i,1}^{C}\left(\hat{\alpha}_{0}+\hat{\alpha}_{1}+\hat{\alpha_{3}}X_{i,1}\right)}{G}+\frac{\sum_{i\in t=0,C}Y_{i,0}^{C}\left(-\hat{\alpha_{0}}-\hat{\alpha_{3}}X_{i,0}\right)}{G}.\end{split}$ (39)
Now, let us interpret the meaning of $\hat{\beta_{3}}$ in Equation (39). To simplify, consider the case where $X_{i,t}$ is a discrete variable and assign it a specific value, say $x$. When $x=0$, Equation (39) reduces to the same form as Equation (14), where $\hat{\beta_{3}}$ represents the “difference-in-difference” of four unconditional means. When $x\neq 0$, Equation (39) can be expressed as follows:
$\begin{split}\hat{\beta}_{3}=&\frac{\sum_{i\in t=1,T}Y_{i,1}^{T}\left(1-\hat{\alpha}_{0}-\hat{\alpha}_{1}-\hat{\alpha}_{2}-\hat{\alpha_{3}}x\right)}{G^{\prime}}-\frac{\sum_{i\in t=0,T}Y_{i,0}^{T}\left(\hat{\alpha}_{0}+\hat{\alpha}_{2}+\hat{\alpha_{3}}x\right)}{G^{\prime}}\\ &-\frac{\sum_{i\in t=1,C}Y_{i,1}^{C}\left(\hat{\alpha}_{0}+\hat{\alpha}_{1}+\hat{\alpha_{3}}x\right)}{G^{\prime}}+\frac{\sum_{i\in t=0,C}Y_{i,0}^{C}\left(-\hat{\alpha_{0}}-\hat{\alpha_{3}}x\right)}{G^{\prime}},\end{split}$ (40)
@@
$\begin{split}G^{\prime}=&\sum_{i\in t=0,T}\left(-\hat{\alpha}_{0}-\hat{\alpha}_{2}-\hat{\alpha_{3}}x\right)^{2}+\sum_{i\in t=1,T}\left(1-\hat{\alpha}_{0}-\hat{\alpha}_{1}-\hat{\alpha}_{2}-\hat{\alpha_{3}}x\right)^{2}\\ &+\sum_{i\in t=0,C}\left(-\hat{\alpha}_{0}-\hat{\alpha_{3}}x\right)^{2}+\sum_{i\in t=1,C}\left(-\hat{\alpha}_{0}-\hat{\alpha}_{1}-\hat{\alpha_{3}}x\right)^{2}.\end{split}$
Continuing with the reasoning from the proof outlined in Appendix A, we can show that $\frac{\left(1-\hat{\alpha}_{0}-\hat{\alpha}_{1}-\hat{\alpha}_{2}-\hat{\alpha}_{3}x\right)}{G^{\prime}}$ and $\frac{\left(\hat{\alpha}_{0}+\hat{\alpha}_{2}+\hat{\alpha}_{3}x\right)}{G^{\prime}}$ are equivalent to the reciprocals of the number of observations in the treated group in the post-intervention period and the pre-intervention period, respectively, with $X_{i,t}=x$. In simpler terms, these represent the probabilities of the observation being in the treated group in their respective time periods, conditional on $X_{i,t}$. It is important to note that this equivalence holds under the assumption of a sufficiently large sample size.
With smaller sample sizes, where the law of large numbers does not apply, therefore, $\hat{\alpha}_{3}\neq\alpha^{0}_{3}$, the true effect of $X_{i,t}$ on the interaction term. Similarly, $\frac{\left(\hat{\alpha}_{0}+\hat{\alpha}_{1}+\hat{\alpha}_{3}x\right)}{G^{\prime}}$ and $\frac{\left(-\hat{\alpha}_{0}-\hat{\alpha}_{3}x\right)}{G^{\prime}}$ represent the probabilities for the observations being in the control group in the post- and pre-intervention periods, respectively, conditional on $X_{i,t}$. As a result, $\hat{\beta_{3}}$ can be expressed as
$\begin{split}&\sum_{i\in t=1,T}P(D_{i}=1,P_{t}=1,X^{T}_{i,1}=x)Y^{T}_{i,1}-\sum_{i\in t=0,T}P(D_{i}=1,P_{t}=0,X^{T}_{i,0}=x)Y^{T}_{i,0}\\ &-\sum_{i\in t=1,C}P(D_{i}=0,P_{t}=1,X^{C}_{i,1}=x)Y^{C}_{i,1}+\sum_{i\in t=0,C}P(D_{i}=0,P_{t}=0,X^{C}_{i,0}=x)Y^{C}_{i,0}.\end{split}$ (41)
@@
$\begin{split}\hat{\beta_{3}}=\left[E[Y^{T}_{i,1}|X^{T}_{i,t}=x]-E[Y^{T}_{i,0}|X^{T}_{i,0}=x]\right]-\left[E[Y^{C}_{i,1}|X^{C}_{i,t}=x]-E[Y^{C}_{i,0}|X^{C}_{i,0}=x]\right].\end{split}$ (42)
This equivalence does not hold when we have small sample sizes. In Equation (42), $\hat{\beta_{3}}$ is the ATT conditional on covariates shown in Equation (29). If $X_{i,t}$ is a continuous random variable, Equation (41) can be rewritten as
@@
UN–DID Regression
When conditional parallel trends is plausible, we can extend the UN–DID regressions shown in equations (6) and (7) to include covariates:
@@
Following the proof shown in Section 3.1, we know that $\hat{\lambda_{2}^{j}}-\hat{\lambda_{1}^{j}}$ from Equation (44) is equivalent to $\hat{\gamma_{1}^{j}}$ from Equation (45). Moving forward, we proceed to derive the coefficient $\hat{\gamma_{1}^{j}}$. Using the FWL theorem, it follows that $\widehat{\gamma_{1}^{j}}$ is equal to the coefficient of $(post^{j}_{t}-\widehat{post^{j}_{t}})$ in the regression presented in Equation (46). Here, $(post^{j}_{t}-\widehat{post^{j}_{t}})$ are the residuals from the model shown in Equation (47):

### v2

@@ 3.3 Results: No Covariates
The simulation study demonstrates that the UN-DID estimator is unbiased. To assess unbiasedness, we generate kernel density plots for both conventional DID and UN-DID estimates of the ATT on a shared axis. The kernel densities for unequal sample sizes are shown in Figure 2. In Panel (a), the kernel density estimates for Case 1 are presented, and in Panel (b), the kernel density estimates for Case 2 are presented. Notably, for Case 1, both kernels are centered around 0, aligning with the true value of the ATT. Similarly, for Case 2, both kernels are centered around 0.1, which is the true value of the ATT for Case 2. The blue line representing the conventional ATT lies exactly beneath the red line representing the UN-DID ATT (except at the tails) and is therefore not visible. This demonstrates that the ATTs of the two methods are exactly numerically equivalent. This finding demonstrates that both conventional and UN-DID estimation methods exhibit unbiasedness, as evidenced by the distributions of estimates being centered around the true values of the ATT. A similar pattern is found with equal sample sizes, in a figure not included.
In order to assess whether the UN-DID approaches the true value as sample size increases, we repeat the Monte Carlo simulation by increasing the sample size to 1,000, 2,000, 4,000, 8,000, 10,000, and 50,000 while maintaining equal sample sizes between both silos. Subsequently, we compute the mean squared error (MSE) between the estimated ATT using UN-DID and the true ATT from the underlying DGP according to Equation (26):
$MSE(\widehat{ATT})=\frac{1}{1000}\sum_{j}^{1000}\biggl{(}\widehat{ATT_{j}}^{UN-DID}-ATT_{0}\biggr{)}^{2}.$ (26)
@@
$MSE(\widehat{SE})=\frac{1}{1000}\sum_{j}^{1000}\biggl{(}\widehat{SE}(\widehat{ATT_{j}})^{UN-DID}-SE_{0}\biggr{)}^{2}.$ (28)
The findings are shown in Figure 3, where the y-axis represents the mean squared error (MSE) and the x-axis corresponds to the sample size. The graph illustrates a decreasing trend in MSE for both the estimated ATT and the standard error as the sample size increases. The results for unequal sample sizes are illustrated in Figure 4. We observe that the MSE for the case with unequal sample sizes is higher than for the case with equal sample sizes but still converges to 0 as the sample size increases.
Figure 3: No Covariates with Equal Sample Sizes: Mean Squared Errors of UN-DID vs. true values
Figure 4: No Covariates with Unequal Sample Sizes: Mean Squared Errors of UN-DID vs. true values
4 UN-DID When Strong Parallel Trends Do Not Hold
The previous section was based on the assumption of strong parallel trends between the treated and the control silos. We now relax Assumption 2 and introduce the conditional parallel trends assumption, so that parallel trends is plausible only conditional on covariates.
@@ 4 UN-DID When Strong Parallel Trends Do Not Hold
$\begin{split}\biggr{[}E[Y_{i}(0)|D_{i}=1,P_{t}=1,X_{i,t}]-E[Y_{i}(0)|D_{i}=1,P_{t}=0,X_{i,t}]\biggr{]}\\ =\biggr{[}E[Y_{i}(0)|D_{i}=0,P_{t}=1,X_{i,t}]-E[Y_{i}(0)|D_{i}=0,P_{t}=0,X_{i,t}]\biggr{]}.\end{split}$
The conditional parallel trends assumption states that the evolution of outcomes between the treated and control groups are the same in the pre-intervention period, conditional on covariates. Under the conditional parallel trends assumption, we can also express the ATT, shown in Equation (3), as the ATT conditional on covariates:
@@
With covariates, we can extend Assumption 6 such that matrices of the covariates, $X_{T}$ and $X_{C}$ for the treated and the control groups, respectively, can be stacked together into a common matrix $X$:
@@
Of course this stacking is not feasible when data are siloed. With unpoolable data, when we gain access to the data from the treated silo, we do not have any information on the data from the control group, or both $Y_{C}$ and $X_{C}$. Similarly, when we gain access to the data from the untreated silo, we do not have any information on the data for the treated silo, or both $Y_{T}$ and $X_{T}$. Because we do not observe the covariates from other groups, we are unable to use non-parametric methods such as the inverse probability weighting (IPW), outcome regression (OR), and the doubly robust methods (DR–DID) to estimate the ATT (Abadie, 2005; Sant’Anna and Zhao, 2020; Heckman et al., 1997).
Researchers typically incorporate covariates in conventional DID to ensure the plausibility of parallel trends given the covariates, particularly when strong parallel trends do not hold (Heckman et al., 1997; Abadie, 2005). Likewise, we are including only covariates that are required for conditional parallel trends to be plausible. For valid estimation of the ATT with covariates, we require that the covariates are correlated with the outcome but not with the treatment on the basis of the conditional independence assumption.
@@
The literature on DID with covariates has largely focused on scenarios where covariates are time-invariant (Abadie, 2005; Callaway and Sant’Anna, 2021; Caetano et al., 2022; Roth et al., 2022; Sant’Anna and Zhao, 2020). When dealing with time-varying covariates, researchers frequently select a specific value of these covariates during a pre-treatment period, as if it were a time-invariant covariate (Caetano et al., 2022).
We will begin by examining the simplest case, where we introduce a single covariate into both the conventional DID and UN-DID regressions. With a single covariate, the conventional DID regression is expressed as
@@
Likewise, we adjust the UN-DID regression to incorporate covariates:
@@
Including time-invariant covariates does not complicate DID analysis, as demonstrated in the following subsection. However, with time-varying covariates, Caetano et al. (2022) has demonstrated that the two-way fixed effects (TWFE) DID regression can provide biased estimate of the ATT due to negative weighting and weight-reversal issues. In this paper, we recognize the complexity introduced by the heterogeneity in data distribution between silos, which can present challenges when analyzing siloed data (Li et al., 2022).
4.1 Common Causal Covariates
One important distinction between the conventional DID regression using poolable data in Equation (31) and the UN-DID regresssions using siloed data in Equations (32) and (33) is how the covariates are handled. In the conventional DID regression there is a single coefficient, $\beta_{4}$, whereas in the UN-DID regressions there are two $\lambda_{3}^{T}$ and $\lambda_{3}^{C}$. Whether or not this matters depends on the relationship between the covariate(s) and the outcome variable and if that varies across silos. Work in progress by Karim and Webb explores this issue in detail. That paper introduces the assumption of common causal covariates, which implies that the relationship between the covariate(s) and the outcome is the same in each silo. The simulations in this paper maintain that assumption. When that assumption is violated, the conventional DID estimator and the estimator from Callaway and Sant’Anna (2021) can be quite biased. Conversely, the UN-DID estimator, or the estimator using pooled data but with silo specific coefficients for covariates introduced by Karim and Webb is unbiased in these settings.
4.2 Equivalence of the Estimated ATT Between Conventional DID and UN-DID Methods with Covariates
Conventional DID Regression
For the conventional DID regression, we assume that Assumption 1 and Assumptions 3 to 7 hold. With a single covariate, the conventional regression is shown in Equation (31). In Equation (31), $\hat{\beta_{3}}$ is the parameter of interest, which is the estimate of the ATT conditional on covariates, and is the sample analogue of Equation (29). According to the FWL theorem, $\hat{\beta_{3}}$ will be the same as the coefficient of $(P_{t}*D_{i}-\widehat{P_{t}*D_{i}})$ from the regression shown in Equation (35). Here, $(P_{t}*D_{i}-\widehat{P_{t}*D_{i}})$ are the residuals from Equation (34):
@@ 4.2 Equivalence of the Estimated ATT Between Conventional DID and UN-DID Methods with Covariates
From Equation (34), we calculate residuals ($P_{t}*D_{i}-\widehat{P_{t}*D_{i}}$) for each observation based on their treatment status and time period. For treated observations in the pre-intervention period, the residual is given by $-\widehat{\alpha}_{0}-\widehat{\alpha}_{2}-\widehat{\alpha}_{3}X_{i,0}$, while for the post-intervention period, it is $1-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{2}-\widehat{\alpha}_{3}X_{i,1}$. Similarly, untreated observations have residuals of $-\widehat{\alpha}_{0}-\widehat{\alpha}_{3}X_{i,0}$ in the pre-intervention period and $-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{3}X_{i,1}$ in the post-intervention period. Subsequently, $\widehat{\beta}_{3}$ from Equation (35) can be estimated using the following OLS formula:
$\begin{split}\widehat{\beta}_{3}=&\sum_{i}\sum_{t}\frac{Y_{i,t}\left(P_{t}*D_{i}-\widehat{P_{t}*D_{i}}\right)}{\left(P_{t}*D_{i}-\widehat{P_{t}*D_{i}}\right)^{2}}\end{split}$ (36)
@@
$\begin{split}&\sum_{i\in t=0,T}Y_{i,0}^{T}\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{2}-\widehat{\alpha}_{3}X_{i,0}\right)+\sum_{i\in t=1,T}Y_{i,1}^{T}\left(1-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{2}-\widehat{\alpha}_{3}X_{i,1}\right)+\\ &\sum_{i\in t=0,C}Y_{i,1}^{C}\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{3}X_{i,0}\right)+\sum_{i\in t=1,C}Y^{C}_{i,1}\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{3}X_{i,1}\right).\end{split}$
@@
$\begin{split}G=&\sum_{i\in t=0,T}\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{2}-\widehat{\alpha}_{3}X_{i,0}\right)^{2}+\sum_{i\in t=1,T}\left(1-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{2}-\widehat{\alpha}_{3}X_{i,1}\right)^{2}\\ &+\sum_{i\in t=0,C}\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{3}X_{i,0}\right)^{2}+\sum_{i\in t=1,C}\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{3}X_{i,1}\right)^{2}.\end{split}$
Combining the numerator and the denominator, $\widehat{\beta}_{3}$ can be written as:
$\begin{split}\widehat{\beta}_{3}=&\frac{\sum_{i\in t=1,T}Y_{i,1}^{T}\left(1-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{2}-\widehat{\alpha}_{3}X_{i,1}\right)}{G}-\frac{\sum_{i\in t=0,T}Y_{i,0}^{T}\left(\widehat{\alpha}_{0}+\widehat{\alpha}_{2}+\widehat{\alpha}_{3}X_{i,0}\right)}{G}\\ &-\frac{\sum_{i\in t=1,C}Y_{i,1}^{C}\left(\widehat{\alpha}_{0}+\widehat{\alpha}_{1}+\widehat{\alpha}_{3}X_{i,1}\right)}{G}+\frac{\sum_{i\in t=0,C}Y_{i,0}^{C}\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{3}X_{i,0}\right)}{G}.\end{split}$ (37)
Now, let us interpret the meaning of $\widehat{\beta}_{3}$ in Equation (37). To simplify, consider the case where $X_{i,t}$ is a discrete variable and assign it a specific value, say $x$. When $x=0$, Equation (37) reduces to the same form as Equation (14), where $\widehat{\beta_{3}}$ represents the “difference-in-difference” of four unconditional means. When $x\neq 0$, Equation (37) can be expressed as follows:
$\begin{split}\widehat{\beta}_{3}=&\frac{\sum_{i\in t=1,T}Y_{i,1}^{T}\left(1-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{2}-\widehat{\alpha}_{3}x\right)}{G^{\prime}}-\frac{\sum_{i\in t=0,T}Y_{i,0}^{T}\left(\widehat{\alpha}_{0}+\widehat{\alpha}_{2}+\widehat{\alpha}_{3}x\right)}{G^{\prime}}\\ &-\frac{\sum_{i\in t=1,C}Y_{i,1}^{C}\left(\widehat{\alpha}_{0}+\widehat{\alpha}_{1}+\widehat{\alpha}_{3}x\right)}{G^{\prime}}+\frac{\sum_{i\in t=0,C}Y_{i,0}^{C}\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{3}x\right)}{G^{\prime}},\end{split}$ (38)
@@
$\begin{split}G^{\prime}=&\sum_{i\in t=0,T}\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{2}-\widehat{\alpha}_{3}x\right)^{2}+\sum_{i\in t=1,T}\left(1-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{2}-\widehat{\alpha}_{3}x\right)^{2}\\ &+\sum_{i\in t=0,C}\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{3}x\right)^{2}+\sum_{i\in t=1,C}\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{3}x\right)^{2}.\end{split}$
Continuing with the reasoning from the proof outlined in Appendix A, we can show that $\frac{\left(1-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{2}-\widehat{\alpha}_{3}x\right)}{G^{\prime}}$ and $\frac{\left(\widehat{\alpha}_{0}+\widehat{\alpha}_{2}+\widehat{\alpha}_{3}x\right)}{G^{\prime}}$ are equivalent to the reciprocals of the number of observations in the treated group in the post-intervention period and the pre-intervention period, respectively, with $X_{i,t}=x$. In simpler terms, these represent the probabilities of the observation being in the treated group in their respective time periods, conditional on $X_{i,t}$. It is important to note that this equivalence holds with a sufficiently large sample size.
With smaller sample sizes, where the law of large numbers does not apply, therefore, $\widehat{\alpha}_{3}\neq\alpha^{0}_{3}$, the true effect of $X_{i,t}$ on the interaction term. Similarly, $\frac{\left(\widehat{\alpha}_{0}+\widehat{\alpha}_{1}+\widehat{\alpha}_{3}x\right)}{G^{\prime}}$ and $\frac{\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{3}x\right)}{G^{\prime}}$ represent the probabilities for the observations being in the control group in the post- and pre-intervention periods, respectively, conditional on $X_{i,t}$. As a result, $\hat{\beta_{3}}$ can be expressed as
$\begin{split}&\sum_{i\in t=1,T}P(D_{i}=1,P_{t}=1|X^{T}_{i,1}=x)Y^{T}_{i,1}-\sum_{i\in t=0,T}P(D_{i}=1,P_{t}=0|X^{T}_{i,0}=x)Y^{T}_{i,0}\\ &-\sum_{i\in t=1,C}P(D_{i}=0,P_{t}=1|X^{C}_{i,1}=x)Y^{C}_{i,1}+\sum_{i\in t=0,C}P(D_{i}=0,P_{t}=0|X^{C}_{i,0}=x)Y^{C}_{i,0}.\end{split}$ (39)
@@
$\begin{split}\widehat{\beta_{3}}=\left[E[Y^{T}_{i,1}|X^{T}_{i,t}=x]-E[Y^{T}_{i,0}|X^{T}_{i,0}=x]\right]-\left[E[Y^{C}_{i,1}|X^{C}_{i,t}=x]-E[Y^{C}_{i,0}|X^{C}_{i,0}=x]\right].\end{split}$ (40)
This equivalence does not hold when we have small sample sizes. In Equation (40), $\widehat{\beta}_{3}$ is the ATT conditional on covariates shown in Equation (29). If $X_{i,t}$ is a continuous random variable, Equation (39) can be rewritten as
@@
UN-DID Regression
When conditional parallel trends is plausible, we can extend the UN-DID regressions shown in equations (6) and (7) to include covariates:
@@
Following the proof shown in Section 3.1, we know that $\widehat{\lambda}_{2}^{j}-\widehat{\lambda}_{1}^{j}$ from Equation (42) is equivalent to $\widehat{\gamma}_{1}^{j}$ from Equation (43). We proceed to derive the coefficient $\widehat{\gamma}_{1}^{j}$. Using the FWL theorem, it follows that $\widehat{\gamma}_{1}^{j}$ is equal to the coefficient of $(post^{j}_{t}-\widehat{post}^{j}_{t})$ in the regression presented in Equation (44). Here, $(post^{j}_{t}-\widehat{post}^{j}_{t})$ are the residuals from the model shown in Equation (45):